# ArticulatePro: A Comparative Study on a Proactive and Non-Proactive Assistant in a Climate Data Exploration Task

### v1

@@ Abstract.
Recent advances in Natural Language Interfaces (NLIs) and Large Language Models (LLMs) have transformed our approach to NLP tasks, allowing us to focus more on a Pragmatics-based approach. This shift enables more natural interactions between humans and voice assistants, which have been challenging to achieve. Pragmatics describes how users often talk out of turn, interrupt each other, or provide relevant information without being explicitly asked (maxim of quantity). To explore this, we developed a digital assistant that constantly listens to conversations and proactively generates relevant visualizations during data exploration tasks. In a within-subject study, participants interacted with both proactive and non-proactive versions of a voice assistant while exploring the Hawaii Climate Data Portal (HCDP). Results suggest that the proactive assistant enhanced user engagement and facilitated quicker insights. Our study highlights the potential of Pragmatic, proactive AI in NLIs and identifies key challenges in its implementation, offering insights for future research.
@@ 1. Introduction
Advances in technology, such as Large Language Models (LLMs) like GPT (Achiam et al., 2023), BERT (Kenton and Toutanova, 2019), and Llama (Touvron et al., 2023), have enabled breakthroughs in various complex Natural Language Processing (NLP) applications. While researchers and policy makers are still resolving challenges regarding the ethics of these powerful innovations, these general-purpose models have made it easier to tackle NLP tasks such as question answering, generating text in the style of deceased authors, and zero-shot or few-shot learning classification (Abbasiantaeb et al., 2024; Pataranutaporn et al., 2023; Bucher and Martini, 2024). As a result, natural language research can now move beyond understanding the semantics and syntax of text (like in text summarization and classification) and shift towards focusing on Pragmatics—understanding the nuances of natural language communication (such as text generation, natural language reasoning, and linguistic presuppositions) (Li et al., 2021). This gives us an opportunity to re-evaluate the way we approach interacting with voice assistants, in order to develop Pragmatic-focused Natural Language applications. For instance, a Pragmatic voice assistant could potentially understand the full range of how people naturally communicate, paving the way for innovations such as AI psychiatrists, elderly caretakers, and personal tutors.
In human conversations, proactivity involves natural interactions—interrupting, talking out of turn, and offering just the right amount of information (Balaraman and Magnini, 2020; Semmens et al., 2019). Current home assistants, like Siri and Alexa, only respond to direct requests, and even advanced AI chatbots like ChatGPT only act when explicitly instructed. This results in AI mirroring the user’s thoughts rather than prompting new ideas or challenging their thinking, which can limit the depth of the conversation (Wilson and Sperber, 1986). Prior research highlights the potential benefits and desire for proactive AI (Völkel et al., 2021; Meurisch et al., 2020; Van Berkel et al., 2021).
To explore the concept of a proactive AI assistant, we focused on the context of data exploration. Over the past decade, Natural Language Interfaces (NLIs) have made significant progress in helping users with data exploration tasks (Shen et al., 2022). These interfaces simplify the process of creating visualizations, allowing users to understand data without programming skills or expertise in visualization (Sun et al., 2010; Quadri et al., 2024). Data exploration serves as our testbed for evaluating the effectiveness of our proactive digital assistant.
@@
To explore these questions, we developed ArticulatePro, a Pragmatic data visualization application, featuring a digital assistant that is capable of proactively generating charts during data exploration tasks. The digital assistant continuously listens to conversations and detects opportunities to generate relevant charts at the right moment. We evaluated our approach through a within-subject user study, where dyads of participants interacted with both a proactive digital assistant and a non-proactive version of the assistant. By comparing participants’ interaction in both conditions, our study aims to uncover potential benefits and challenges of a proactive digital assistant. In our study, participants explored the Hawaii Climate Data Portal (HCDP), a climate data repository developed by the University of Hawaii at Manoa (McLean et al., 2020, 2023; Mclean and Cleveland, 2023; Longman et al., 2024).
@@
    Determined the effectiveness of our proactive digital assistant.
@@
    We compare the interactions between a proactive digital assistant and a non-proactive digital assistant.
@@
    We highlight the usability challenges encountered when working with a proactive digital assistant for data exploration tasks.
@@ 2.1. Natural Language for Data Exploration
In the past decade, we have made significant improvements towards the research of systems that facilitate data exploration tasks that utilize natural language voice commands. This focus allows users to conduct data exploration tasks more efficiently, directing their attention to exploration rather than chart generation. In our prior work, we developed and evaluated Articulate (Sun et al., 2010), one of the earliest natural language-enabled data exploration systems (Shen et al., 2022). We found that users took, on average, 12 times faster to create a chart than using Excel to create a chart. This performance improvement introduces the efficiency that can be leveraged using a natural language system.
@@
In 2015 and 2016, Gao et al and Setlur et al focused on providing user interface elements for users to manipulate their verbally-created charts (Gao et al., 2015; Setlur et al., 2016). These studies focused on a hybrid interaction, where users can still utilize the keyboard and mouse while also using text interaction. This paper focuses primarily on voice commands for chart construction rather than relying on other modalities.
@@
In 2016, we conducted a Wizard-of-Oz study, revealing that 15% of user utterances were queries, while the remaining 85% provided context for queries during a data exploration task (Kumar et al., 2016). Queries lacking sufficient context failed to generate the user’s intended visualizations. From 2018 to 2023, Hoque et al., Srinivasan et al., Setlur et al., and our prior work, designed context-supported data exploration systems (Setlur et al., 2016; Srinivasan and Setlur, 2021; Hoque et al., 2017; Tabalba et al., 2022, 2023b). We have all shown how context repairs queries that were imprecise or incomplete. These studies highlight the need for more context-supported natural language systems, one of the key aspects of Pragmatics.
@@
From 2017 through 2023, we iterated on a conversational data exploration system, focusing our efforts on resolving co-reference resolution (Aurisano et al., 2015, 2016; Kumar et al., 2017; Bhattacharya et al., 2023). LLMs have been shown to perform exceptionally well at addressing co-reference resolution (Gan et al., 2024), which is one example where LLM’s generalized knowledge can be leveraged to address previously difficult NLP-problems.
2.2. Proactive Digital Assistants
@@
The need for proactive systems has been discussed in the literature. In one study, McMillan et al. hired Professional designers to analyze recorded daily conversations of users (McMillan et al., 2015). The professional designers proposed potential applications that utilize conversational data. One of the applications the designers suggested was proactively launching mobile applications. For instance, if keywords like “hungry” and “food” were detected by the mobile device, it could launch Uber Eats, a popular food delivery application. Volkel et al., surveyed 205 participants on how they envision the “perfect voice assistant” (Völkel et al., 2021). They found “Majority of participants let their voice assistant take the lead in parts of the dialogue”. Specifically, 87% of participants wanted the voice assistant to make suggestions, 83% wanted it to be an automatic scheduler, and 80% advocated for assistant-led dialogue. These findings suggest that users desire a proactive voice assistant.
@@
Reicherts et al., conducted an online survey to explore user perceptions on storyboards of proactive voice assistants (Reicherts et al., 2021). Participants perceived the proactiveness of the voice assistant to be useful, appropriate, and pleasant. Zargham et al., conducted a similar online study evaluating user perceptions on usefulness, appropriateness, and invasiveness (Zargham et al., 2022). They introduced the term “The Proactivity Dilemma” to describe the dual nature of NLIs initiating conversations, acting as both beneficial and a potential intrusion, depending on the context. This work illuminates how possible opportunities for proactivity affect the user’s perceptions on its intrusiveness. In our paper, we study how users work with a working proactive assistant, rather than evaluating storyboards. We also only focus on a single opportunity of proactivity, comparing its effectiveness against a non-proactive assistant in a data exploration task.
@@
Andolina et al. developed a system that could proactively retrieve Google search results, allowing users to focus on interpreting the results rather than explicitly formulating their query (Andolina et al., 2018). Our study differs as we focus on the context of data exploration tasks with dyads of participants rather than the context of retrieving Google search results with individual participants. Shi et al. developed a brainstorming system that could retrieve relevant images representing user conversations (Shi et al., 2017). Their study revealed how their constantly listening brainstorming system can help users produce more ideas and reduce the number of lulls. Shi et al. focus on evaluating the effectiveness of displaying images while we compare the differences of how users work between a proactive assistant and a non-proactive assistant in a data exploration task. Balaraman and Bernardo focus on building a proactive information retrieval system (Balaraman and Magnini, 2020), called SimDial. Similar to our approach, Simdial also focuses on Pragmatics, but differentiates itself by evaluating their approach on short simulated dialogues. They describe how their proactive approach can reduce the number of back-and-forth dialogues between the user and computer. In our paper, we implement a working proactive digital assistant and conduct a data exploration study to evaluate its effectiveness, as opposed to evaluating simulated dialogues.
@@ 2.3. Pragmatics
Li et al. offer a comprehensive overview of Natural Language Processing (NLP), highlighting three pivotal areas: semantics, syntax, and pragmatics (Li et al., 2021). Their synthesis indicates that while substantial progress has been made in semantics and syntax, there is a pressing need to shift NLP research toward a pragmatic approach, emphasizing the incorporation of pragmatics in the development of Natural Language Interfaces (NLIs).
Pragmatics, a branch of linguistics, delves into the nuances of natural language conversations. A key aspect of pragmatics is context. According to Cutting, context encompasses the physical location of the conversation, the background knowledge generally known, and the knowledge shared between participants (Cutting, 2005). Depending on context, the entire meaning of a sentence can be changed (Grice, 1975). Paul Grice, a prominent figure in pragmatics, introduced the concept of ”implicatures,” which extend beyond the semantic and syntactic dimensions of verbal or written communication. In our system’s development, we incorporate context by allowing the system to continuously listen to the conversation rather than solely processing the user’s query.
Grice also proposed the Cooperative Principles, which are observations of rules that people instinctively follow during conversations. These principles include quantity, quality, relevance, and manner. The maxim of quantity suggests providing just enough information, while the maxim of quality emphasizes delivering information believed to be true. The maxim of relevance dictates that communication should be pertinent to the current discussion, and the maxim of manner advises that communication should be clear and unambiguous. It is important to note that these principles are based on interactions between cooperating individuals. In real conversations, people may break these maxims for various reasons.
Sperber and Wilson further developed Grice’s ideas, positing that all conversations can be reduced to the maxim of relevance (Wilson and Sperber, 1986). They argue that this principle simplifies communication by focusing on what is relevant to establish common ground between conversing individuals. This common ground, known as the ‘cognitive environment,’ encompasses the shared knowledge between participants. In our system, we aim to apply Sperber and Wilson’s theory of relevance to our pragmatic approach, ensuring that our system can effectively engage in meaningful and contextually appropriate interactions.
@@ 3.1. An Overview of the System
ArticulatePro is implemented as an application in SAGE3 (Smart Amplified Group Environment), a web app for managing large tiled display walls (Tabalba et al., 2023a; Harden et al., 2023). SAGE3 is developed using Javascript as its frontend and Node and Python as its backend. We chose SAGE3 as the substrate for ArticulatePro as it allowed us to take advantage of its features, such as its capability to manage content on an infinite canvas, while simplifying the development stack needed to create an application.
Figure 2. (1) General Workspace to move, resize and delete selected visualizations. (2) Digital Persona of ArticulatePro that creates visualizations. (3) Visualization Conveyor belt (inspired by Sushi Conveyor belt) displays generated visualizations from the digital assistant. The digital assistant creates visualizations and displays them in the conveyor belt. Users select charts from the conveyor belt where they are moved to the general workspace for further analysis.
The ArticulatePro application, in Figure 2, is a visualization application where users can interact with a digital assistant, Arti, that can assist in data exploration tasks through natural language voice commands. In our Pragmatic approach, we designed Arti to take advantage of conversational context by continuously listening to the conversation. Then, based on what Arti hears, it will generate relevant charts. Arti determines the best time to generate visualizations, either through an explicit request or proactively, similar to how a human might respond in a conversation. Here are short descriptions on what an Explicit Request is and how Arti proactively creates charts:
@@
When the user explicitly asks for a chart, Arti behaves similarly to current digital assistants like Siri, Alexa, and Cortana, but does not require wake words. Arti listens continuously and generates a chart whenever it detects a command. As an example, the user might say “Generate me a chart on Covid Risk and Diabetes Rate”. In this example, the user is directly asking the assistant to generate a chart showing two variables.
Proactively:
These are instances where Arti tries to be proactive when the user isn’t explicitly asking for a chart. For example, if two users are exploring medical data and one says, ”I don’t get why COVID rates are higher in the East than in the West,” Arti might generate a chart comparing medical factors between the East and West, even though the user didn’t directly ask for it. This could help clarify the user’s confusion by providing additional context on the discrepancy of COVID rates in the East and West. This example shows how Arti can be proactive in data exploration tasks.
Once Arti decides to generate a chart, either by explicit request or proactively, it will display the chart on the visualization conveyor belt (Figure 2, Box 3). Users can hover over the chart to see a larger preview in the center of the screen for easier inspection. If they find the chart valuable, they can click on it to move it to the workspace (Figure 2, Box 1), where it can be rearranged, resized, compared with other charts, or deleted.
In the next sections, we describe how we designed our proactive approach. First we explain our investigation of exploring opportunities for a digital assistant to be proactive in a data exploration task. Then we describe our proactive solution by leveraging Pragmatic theories of how people communicate.

### v2

@@ Abstract.
Recent advances in Natural Language Interfaces (NLIs) and Large Language Models (LLMs) have transformed our approach to NLP tasks, shifting the focus towards a more Pragmatics-based approach. This shift enables more natural interactions between humans and voice assistants, which have been historically difficult to achieve. Pragmatics involves understanding how users often talk out of turn, interrupt one another, or provide relevant information without being explicitly asked (maxim of quantity). To explore this, we developed a digital assistant that continuously listens to conversations and proactively generates relevant visualizations during data exploration tasks. In a within-subject study, participants interacted with both proactive and non-proactive versions of a voice assistant while exploring the Hawaii Climate Data Portal (HCDP). Results suggest that the proactive assistant enhanced user engagement and facilitated quicker insights. Our study highlights the potential of Pragmatic, proactive AI in NLIs and identifies key challenges in its implementation, offering insights for future research.
@@ 1. Introduction
Advances in technology, such as Large Language Models (LLMs) like GPT (Achiam et al., 2023), BERT (Kenton and Toutanova, 2019), and Llama (Touvron et al., 2023), have enabled breakthroughs in various complex Natural Language Processing (NLP) applications. While researchers and policymakers are still addressing the ethical challenges posed by these powerful innovations, these general-purpose models have made it easier to tackle NLP tasks such as question answering, generating text in the style of deceased authors, and zero-shot or few-shot learning classification (Abbasiantaeb et al., 2024; Pataranutaporn et al., 2023; Bucher and Martini, 2024). As a result, natural language research can now move beyond the focus on understanding semantics and syntax (as seen in text summarization and classification) and shift towards Pragmatics—understanding the nuances of natural language communication, such as text generation, natural language reasoning, and linguistic presuppositions (Li et al., 2021). This shift gives us the opportunity to re-evaluate how we approach interacting with voice assistants to develop Pragmatics-focused Natural Language applications. For instance, a Pragmatic voice assistant could potentially understand the full range of how people naturally communicate, paving the way for innovations such as AI psychiatrists, elderly caretakers, and personal tutors.
In human conversations, proactivity involves natural interactions—interrupting, talking out of turn, and offering just the right amount of information (Balaraman and Magnini, 2020; Semmens et al., 2019). Current home assistants, like Siri and Alexa, only respond to direct requests, and even advanced AI chatbots like ChatGPT act only when explicitly instructed. This results in AI mirroring the user’s thoughts rather than prompting new ideas or challenging their thinking, which can limit the depth of conversations (Wilson and Sperber, 1986). Prior research highlights the potential benefits and the growing demand for proactive AI (Völkel et al., 2021; Meurisch et al., 2020; Van Berkel et al., 2021).
To explore the concept of a proactive AI assistant, we focused on the context of data exploration. Over the past decade, Natural Language Interfaces (NLIs) have made significant progress in assisting users with data exploration tasks (Shen et al., 2022). These interfaces simplify the process of creating visualizations, enabling users to understand data without programming skills or expertise in visualization (Sun et al., 2010; Quadri et al., 2024). Data exploration serves as our testbed for evaluating the effectiveness of our proactive digital assistant.
@@
To explore these questions, we developed ArticulatePro, a Pragmatic data visualization application featuring a digital assistant capable of proactively generating charts during data exploration tasks. The digital assistant continuously listens to conversations and detects opportunities to generate relevant charts at the right moment. We evaluated our approach through a within-subject user study, where dyads of participants interacted with both a proactive digital assistant and a non-proactive version of the assistant. By comparing participants’ interactions in both conditions, our study aims to uncover the potential benefits and challenges of a proactive digital assistant. In our study, participants explored the Hawaii Climate Data Portal (HCDP), a climate data repository developed by the University of Hawaii at Manoa (McLean et al., 2020, 2023; Mclean and Cleveland, 2023; Longman et al., 2024).
@@
    An evaluation of the effectiveness of our proactive digital assistant.
@@
    A comparison of interactions between a proactive and a non-proactive digital assistant.
@@
    A discussion of the usability challenges encountered when working with a proactive digital assistant for data exploration tasks.
@@ 2.1. Natural Language for Data Exploration
Over the past decade, significant advancements have been made in the research of systems that facilitate data exploration tasks using natural language voice commands. This focus enables users to conduct data exploration more efficiently, allowing them to concentrate on the exploration process rather than on chart generation. In our prior work, we developed and evaluated Articulate (Sun et al., 2010), one of the earliest natural language-enabled data exploration systems (Shen et al., 2022). We found that users, on average, created charts 12 times faster than when using Excel, showcasing the efficiency gains that can be achieved with a natural language system.
@@
In 2015 and 2016, Gao et al. and Setlur et al. explored user interface elements that allowed users to manipulate their verbally-created charts (Gao et al., 2015; Setlur et al., 2016). These studies introduced hybrid interactions, where users could still utilize the keyboard and mouse alongside text interaction. However, this paper focuses primarily on voice commands for chart creation, rather than relying on other modalities.
@@
In 2016, we conducted a Wizard-of-Oz study, revealing that only 15% of user utterances were direct queries, while the remaining 85% provided context for queries during data exploration tasks (Kumar et al., 2016). Queries lacking sufficient context often failed to generate the intended visualizations. From 2018 to 2023, researchers including Hoque et al., Srinivasan et al., Setlur et al., and our own prior work, designed context-supported data exploration systems (Setlur et al., 2016; Srinivasan and Setlur, 2021; Hoque et al., 2017; Tabalba et al., 2022, 2023b). These studies demonstrated how incorporating context could help repair imprecise or incomplete queries, emphasizing the importance of context-supported natural language systems—one of the key aspects of Pragmatics.
@@
Between 2017 and 2023, we iterated on a conversational data exploration system, concentrating on resolving co-reference resolution (Aurisano et al., 2015, 2016; Kumar et al., 2017; Bhattacharya et al., 2023). Large Language Models (LLMs) have been shown to perform exceptionally well in addressing co-reference resolution (Gan et al., 2024), showcasing how LLMs’ generalized knowledge can be leveraged to solve previously challenging NLP problems.
2.2. Proactive Applications
@@
The need for proactive systems has been widely discussed in the literature. In one study, McMillan et al. hired professional designers to analyze recorded daily conversations of users (McMillan et al., 2015). The designers proposed potential applications that could utilize conversational data. One suggestion was to proactively launch mobile applications. For example, if keywords like ”hungry” or ”food” were detected by the mobile device, it could automatically launch Uber Eats, a popular food delivery application. Volkel et al. surveyed 205 participants on their vision of the ”perfect voice assistant” (Völkel et al., 2021). They found that the majority of participants wanted their voice assistant to take the lead in parts of the dialogue. Specifically, 87
@@
Reicherts et al. conducted an online survey to explore user perceptions of storyboards featuring proactive voice assistants (Reicherts et al., 2021). Participants found the proactiveness of the assistant to be useful, appropriate, and pleasant. Similarly, Zargham et al. conducted an online study evaluating user perceptions of usefulness, appropriateness, and invasiveness (Zargham et al., 2022). They introduced the term ”The Proactivity Dilemma” to describe the dual nature of Natural Language Interfaces (NLIs) initiating conversations—being both beneficial and potentially intrusive, depending on the context. This work highlights how opportunities for proactivity can affect users’ perceptions of its intrusiveness. In our paper, we study how users interact with a working proactive assistant, rather than evaluating storyboards. Additionally, we focus on a specific instance of proactivity, comparing its effectiveness against a non-proactive assistant in a data exploration task.
@@
Andolina et al. developed a system that could proactively retrieve Google search results, allowing users to focus on interpreting the results rather than formulating a query explicitly (Andolina et al., 2018). Our study differs by focusing on data exploration tasks with dyads of participants rather than retrieving Google search results with individual participants. Shi et al. developed a brainstorming system that could retrieve relevant images based on user conversations (Shi et al., 2017). Their study demonstrated how a constantly listening system could help users generate more ideas and reduce the number of conversational lulls. While Shi et al. evaluated the effectiveness of displaying images, we compare how users interact with a proactive assistant versus a non-proactive assistant during data exploration tasks. Balaraman and Bernardo developed a proactive information retrieval system called SimDial (Balaraman and Magnini, 2020), which, like our approach, focuses on Pragmatics but differs by evaluating short simulated dialogues. Their proactive approach reduced the number of back-and-forth exchanges between the user and the system. In our study, we implement a fully functional proactive digital assistant and evaluate its effectiveness in a real-world data exploration setting, as opposed to simulated dialogues.
@@ 2.3. Pragmatics
Li et al. provide a comprehensive overview of Natural Language Processing (NLP), highlighting three pivotal areas: semantics, syntax, and pragmatics (Li et al., 2021). Their analysis indicates that while substantial progress has been made in semantics and syntax, there is an urgent need to shift NLP research toward Pragmatics, emphasizing its incorporation into the development of Natural Language Interfaces (NLIs).
Pragmatics, a branch of linguistics, deals with the nuances of natural language conversations. A key element of Pragmatics is context. According to Cutting, context encompasses the physical location of the conversation, the shared background knowledge, and the mutual understanding between participants (Cutting, 2005). The meaning of a sentence can change entirely depending on the context (Grice, 1975). Paul Grice, a prominent figure in Pragmatics, introduced the concept of ”implicatures,” which extend beyond the semantic and syntactic dimensions of verbal or written communication. In our system, we incorporate context by continuously listening to the conversation, rather than only processing isolated user queries.
Grice also proposed the Cooperative Principles, which are guidelines that people typically follow during conversations. These principles include quantity, quality, relevance, and manner. The maxim of quantity suggests providing just enough information, while the maxim of quality emphasizes delivering information believed to be true. The maxim of relevance dictates that communication should be pertinent to the current discussion, and the maxim of manner advises that communication should be clear and unambiguous. It is important to note that these principles are based on interactions between cooperating individuals, and in real conversations, people may break these maxims for various reasons.
Sperber and Wilson further developed Grice’s ideas, arguing that all conversations can be reduced to the maxim of relevance (Wilson and Sperber, 1986). They posit that this principle simplifies communication by focusing on what is relevant, which helps to establish common ground between conversing individuals. This common ground, referred to as the ”cognitive environment,” includes the shared knowledge between participants. In our system, we aim to apply Sperber and Wilson’s theory of relevance, ensuring that our system engages in meaningful and contextually appropriate interactions.
@@ 3.1. An Overview of the System
ArticulatePro is implemented as an application in SAGE3 (Smart Amplified Group Environment), a web app designed for managing large tiled display walls (Tabalba et al., 2023a; Harden et al., 2023). SAGE3 is developed using JavaScript for the frontend and Node.js and Python for the backend. We chose SAGE3 as the foundation for ArticulatePro because it allowed us to leverage its features, such as managing content on an infinite canvas, while simplifying the development stack needed to build the application.
Figure 2. (1) General workspace for moving, resizing, and deleting selected visualizations. (2) The digital persona of ArticulatePro that creates visualizations. (3) A visualization conveyor belt (inspired by a sushi conveyor belt) that displays generated visualizations from the digital assistant. The assistant creates visualizations and displays them on the conveyor belt, where users can select charts to move them to the general workspace for further analysis.
The ArticulatePro application, shown in Figure 2, is a visualization tool where users can interact with a digital assistant, Arti, to assist in data exploration tasks using natural language voice commands. In our Pragmatic approach, we designed Arti to take advantage of conversational context by continuously listening to user conversations. Based on what Arti hears, it generates relevant charts. Context is a critical aspect of Pragmatics (Cutting, 2005), and by contextualizing user queries, Arti is able to resolve queries that may be incomplete or ambiguous. Arti determines the best time to generate visualizations, either in response to an explicit request or proactively, much like a human might respond in a conversation. Below are descriptions of an explicit request and how Arti proactively creates charts:
@@
When the user explicitly asks for a chart, Arti behaves similarly to current digital assistants like Siri, Alexa, and Cortana, but without the need for wake words. Arti listens continuously and generates a chart whenever it detects a command. For example, the user might say, “Generate a chart on COVID risk and diabetes rate.” In this case, the user directly requests a chart that shows the two variables.
Proactive:
In proactive instances, Arti generates a chart even when the user hasn’t explicitly requested one. For example, if two users are exploring medical data and one says, ”I don’t get why COVID rates are higher in the East than in the West,” Arti might generate a chart comparing medical factors between the East and West, even though the user didn’t directly ask for it. This could help clarify the user’s confusion by providing additional context on the discrepancy in COVID rates between the East and West. This example illustrates how Arti can be proactive during data exploration tasks.
Once Arti generates a chart—either through an explicit request or proactively—it is displayed on the visualization conveyor belt (Figure 2, Box 3). Users can hover over the chart to see a larger preview in the center of the screen for easier inspection. If they find the chart valuable, they can click on it to move it to the workspace (Figure 2, Box 1), where it can be rearranged, resized, compared with other charts, or deleted.
In the following sections, we describe how we designed our proactive approach. First, we explain our investigation into exploring opportunities for a digital assistant to be proactive during data exploration tasks. Then, we describe our proactive solution, leveraging Pragmatic theories of communication.